# FaithLM: Towards Faithful Explanations for Large Language Models

### v3

@@ 3.2 FaithLM for Fidelity-enhanced Explanation
The framework of Fidelity-enhanced Explanation is illustrated in Figure 2(a). Since the initial explanation may be unreliable and unfaithful, we propose a fidelity-enhanced optimization approach designed to progressively generate explanations with higher fidelity. We aim to explain the response $\boldsymbol{Y}$ produced by the targeted LLM $f(\cdot)$ in response to the given input queries $\boldsymbol{X}$ with NL explanations $\mathcal{E}_{\text{NL}}$ following the goal of fidelity enhancement. In the first round of enhancement, the LLM explainer generates NL explanations $\mathcal{E}_{\text{NL}}$ following human-crafted explanation trigger prompts $\mathcal{P}_{\mathcal{E}}$ provided in Appendix G. The explanations are then generated by the explainer $\textsl{g}(\mathcal{P}_{\mathcal{E}}|~{}\boldsymbol{X},\boldsymbol{Y})$. Starting from the second round till converge, FaithLM collects a trajectory $\mathcal{T}$ with the NL explanations $\mathcal{E}_{\text{NL}}$ and their corresponding fidelity scores $\mathcal{S}_{\mathcal{E}}$ generated by Fidelity Evaluator. The collection process can be represented as $\mathcal{T}\leftarrow\{\mathcal{T},~{}[\mathcal{E}_{\text{NL}},~{}\mathcal{S}_{\mathcal{E}}]\}$, where $\mathcal{T}$ initially starts as an empty trajectory. Following this trajectory, the LLM explainer generates new explanations with the goal of achieving higher fidelity scores in subsequent iterations. This fidelity enhancement process is guided by the system prompts detailed in Figure 13 of Appendix H.
@@
(a) Fidelity-enhanced Explanation
(b) Trigger Prompt Optimization
Figure 2: An overview of FaithLM framework for two differenent optimization objectives. The blue dotted line reveals the trajectory to optimize the NL explanation (Section 3.2), and the red dotted line indicates the trajectory of the explanation trigger prompt optimization (Section 3.3). “Traj. Prompt" denotes the trajectory system prompt shown in Section H.
Algorithm 1 FaithLM for fidelity-enhanced explanation
@@
1:  $\mathcal{E}_{\text{NL}}\sim\textsl{g}(\mathcal{P}_{\mathcal{E}}~{}|~{}\boldsymbol{X},\boldsymbol{Y})$
2:  $\mathcal{T}=\varnothing$
@@
Algorithm of Fidelity-enhanced Explanation.
The outline of FaithLM for Fidelity-enhanced Explanation is detailed in Algorithm 1. Specifically, in the first iteration, FaithLM generates the NL explanations using the human-craft prompts (line 1). starting from the second iteration till the convergence or optimization ends, FaithLM estimates the fidelity of the derived NL explanations (line 4). Then, we incorporate the explanation and its corresponding fidelity score to the trajectory (line 5), and update the explanations with the goal of achieving higher fidelity scores in subsequent iterations (line 6). The iteration terminates at a predetermined step or ceases earlier as soon as FaithLM observes a flipping performance from the targeted LLM $f(\cdot)$.
3.3 FaithLM for Trigger Prompt Optimization
Despite the success of enhancing fidelity in Section 3.2, the low quality of the explanation trigger prompts $\mathcal{P}_{\mathcal{E}}$ may still hinder the optimization process of receiving a high-fidelity explanation. Given that the unknown preference for prompts from LLMs, human-crafted trigger prompts used in Fidelity-enhanced Explanation Optimization might lead to sub-optimal fidelity enhancement in the derived explanations. In this section, we hereby propose a new optimization pipeline under FaithLM, aiming to optimize the trigger prompt $\mathcal{P}_{\mathcal{E}}$ for generating explanations with higher fidelity.
@@ 3.3 FaithLM for Trigger Prompt Optimization
The framework of Trigger Prompt Optimization is shown in Figure 2(b). The framework aims to optimize the trigger prompt to generate NL explanations with higher fidelity. Different from the optimization goal in Section 3.2, the trajectory in this task collects the trigger prompts $\mathcal{P}_{\mathcal{E}}$ and their fidelity scores $\mathcal{S}_{\mathcal{P}}$. The trajectory is constructed by the system optimization prompts detailed in Figure 12 of Appendix H.
To estimate the fidelity score for a trigger prompt, FaithLM first adopts the randomly human-crafted trigger prompt to guide the LLM explainers to generate NL explanations, and then utilize the Fidelity Evaluator to assess the fidelity of the derived explanation. The final estimated score is averaged by the fidelity score $\mathcal{S}_{\mathcal{E}_{i}}$ of each instance $i$ from the hold-out dataset $(\boldsymbol{X}_{i},\boldsymbol{Y}_{i})\in\mathcal{D}$. Formally, the fidelity score for a trigger prompt $\mathcal{P}_{\mathcal{E}}$ can be estimated as follows:
@@
where $\mathcal{S}_{\mathcal{E}{i}}$ represents the fidelity score of the explanation $\mathcal{E}_{i}$, which is generated by $\textsl{g}(\mathcal{P}_{\mathcal{E}}|\boldsymbol{X}_{i},\boldsymbol{Y}_{i})$, as assessed by the Fidelity Evaluator discussed in Section 3.1.
During the optimization procedure, the trajectory begins from an empty set and starts to incorporate newly derived trigger prompts along with the fidelity scores in each optimization iteration. Following this trajectory, the LLM explainer generates a new trigger prompt with the goal of achieving higher fidelity scores of explanations in subsequent iterations. After several rounds of iterations, FaithLM ultimately yields an optimal explanation trigger prompt with the highest fidelity score for the LLM explainer to generate a more faithful NL explanation.
@@
The outline of FaithLM for Trigger Prompt Optimization is detailed in Algorithm 2, which focuses on optimizing the trigger prompt for generating NL explanations. Specifically, in each iteration, LLM explainer $\textsl{g}(\cdot)$ leverages the trigger prompt to generate the NL explanations and estimates its fidelity (lines 4-7). The fidelity scores of the trigger prompts average the fidelity scores of the entire hold-out dataset (lines 8). Afterward, the trajectory appends the trigger prompt with its corresponding fidelity score (line 9), and updates the trigger prompt as a new sequence of words to achieve higher fidelity scores (line 10). Through multiple iterations, FaithLM progressively guides the trigger prompt to generate explanations with higher fidelity scores, following the reference path established in the trajectory. The iteration process terminates at a predetermined step, and we set 20 in our experiments.
@@ 4.1 Dataset and Baseline
Datasets. We evaluate FaithLM on three datasets with multiple tasks: ECQA [2] dataset on commonsense question-answer task, TrivaQA-Long [5, 19] dataset on reading comprehension task, and COPA [20, 35] dataset on commonsense causal reasoning task. More details of datasets are provided in Appendix A. Baseline Methods. We compare FaithLM with two state-of-the-art baseline methods: SelfExp [26] and Self-consistency [42]. The former ones instruct LLMs to generate explanations using prompt engineering under single-forward inference, and the later ones leverage the outputs from the chain-of-thought prompting process as the model explanations.
4.2 Experiment Settings
In this part, we introduce the experimental settings for evaluating FaithLM. Two distinct types of explanation tasks and evaluation settings are illustrated as follows.
Fidelity-enhanced Explanation
In this task, our goal is to produce NL explanations that exhibit a higher degree of fidelity. The fidelity is exploited as a metric to evaluate fidelity. FaithLM is evaluated across all testing instances, where an NL explanation is generated for each instance, and the averaged fidelity score is calculated, serving as the reported metric to evaluate fidelity.
Explanation Trigger Prompt Optimization.
Evaluation Metrics.
The quality of the derived NL explanation is evaluated under the fidelity and truthfulness metrics. The fidelity follows Section 3.1, which observes the flipping rate of the targeted LLMs by incorporating contrary explanations to the input. The evaluation of truthfulness assesses the correlation between the derived NL explanations to the ground-truth explanations Specifically, we leverage GPT-3.5 and two well-trained natural language inference (NLI) models, Roberta-Large and XLNet-Large [31] from the huggingface hub [43], as the evaluators. With the same evaluators setup, the truthfulness evaluation follows the settings from [23], and uses the evaluation prompt provided in Appendix F. Specifically, the evaluators assess the derived explanations and ground-truth explanations, determining whether the two sentences belong to “similar content", “dissimilar content," or “non-relevant content". The higher proportion of “similar content", the more consistent results with the ground-truth NL explanations.
Implementation Details.
In the experiments, we explore two variants of LLMs as the targeted LLMs $f(\cdot)$: Vicuna-7B [11] and Phi-2 [18], two types of LLMs as the explainers $\textsl{g}(\cdot)$ in FaithLM: GPT-3.5-Turbo and Claude-2 [3]. The LLM agent for generating the contrary explanations takes the same LLMs as those used by the explainers. All reported results are calculated from the average scores of 3 times repetitions. The hyper-parameter settings and device configuration, including temperature and total optimization steps, of FaithLM are given in Appendix C and D, respectively.
@@ 4.3 Explanation Efficacy of FaithLM (RQ1)
In this section, we explore the efficacy of FaithLM from two perspectives. Specifically, we first evaluate the efficacy of NL explanations generated by Algorithm 1, and then assess the explanation trigger prompts generated by Algorithm 2.
@@
We assess the efficacy of derived explanations under the fidelity metric. FaithLM adopts the trajectory system prompts in Figure 13 of Appendix H. The generation of contrary explanations is guided by the prompt in Table 8 of Appendix G.
Figure 3: The fidelity evaluation of explanations from FaithLM on ECQA dataset (left), TriviaQA-Long (middle), and COPA dataset (right). The reported scores are the average fidelity on testing instances in each step of fidelity-enhanced optimization.
Figure 4: Trustfulness evaluation of the NL explanations. The higher the proportion of “similar to ground-truth explanation" obtained from classifiers, the more consistent the derived explanations are with the ground-truth NL explanations.
  • Fidelity Evaluation. The results in Figure 3 demonstrate that FaithLM achieves significantly higher fidelity scores across all three datasets compared with two baselines after 20 steps of optimization. Moreover, the optimization curve of fidelity demonstrates that 20 rounds of optimization are sufficient to converge. Similar phenomenon occurs across different settings of explainers and targeted LLMs. Additional results are provided in Appendix E.
  • Truthfulness Evaluation. To evaluate the truthfulness of the derived explanations, we show the proportions of “similar to ground-truth explanations" in the ECQA dataset, as depicted in Figure 4. We leverage GPT evaluator and well-trained NLI evaluators to determine whether the derived explanations from FaithLM and ground-truth explanations are within similar content or not. The results show that FaithLM’s derived explanations are more consistent with the ground-truth NL explanations, which is indicated by a larger proportion of “similar to ground-truth explanations" generated by FaithLM than baseline methods.
Efficacy of Explanation Trigger Prompts.
We first show prompt optimization curves of explanation trigger prompts on three different datasets, and then leverage the optimal explanation trigger prompts to generate explanations via FaithLM. In the experiments, we randomly select 15 instances from the training dataset in each optimization round, and compute the average fidelity scores of the newly derived trigger prompts. After the progress is terminated, we evaluate the optimized trigger prompts on the testing set. The optimization step is uniformly established at 50 rounds across different explainer and targeted LLM settings. We evaluate the derived explanation trigger prompts under two distinct experimental settings, as shown below.
  • Trigger Prompt Optimization Curve. Figure 5 demonstrates the optimization curves of three datasets. We display the explainer as GPT-3.5-Turbo and Claude-2 and the explainer as Vicuna-7B. We observe that the optimization curve exhibits a generally ascending trend as the step progresses, interspersed with multiple waves throughout the optimization procedure. This indicates that FaithLM generates better explanation trigger prompts after the optimization. More results of optimization curves on remaining datasets are provided in Appendix E.
  • Explanation Generation by Optimized Trigger Prompts. We utilize the optimized explanation trigger prompts to generate explanations following Algorithm 1. The results are displayed in Figure 6(a), including the experiments conducted using all three datasets with Claude-2 as the explainer and Vicuna-7B as the targeted LLM. We observe that optimized explanation trigger prompts obtain higher fidelity scores than the initial human-crafted trigger prompt in generating explanations. This trend is consistent across all datasets, regardless of whether the explanations are refined by Algorithm 1. Additionally, the optimized trigger prompts are effective in improving the fidelity of the derived explanations with or without the optimization of FaithLM.
A Case Study of  FaithLM.
The case studies illustrate the evolving trend via FaithLM, including derived NL explanations, explanation trigger prompts, and contrary explanations in Appendix I. These studies demonstrate that the explanations generated by FaithLM are informative and readable, which enables humans to understand the reasons behind the decision-making process of LLMs.
4.4 Transferability of Optimized Trigger Prompt (RQ2)
We assess the transferability of ultimately optimized trigger prompts across different unseen datasets within the same domain, as depicted in Figure 6(b). Specifically, we transfer the optimized trigger prompts from the ECQA to the Social-IQA dataset, and from the COPA to the XCOPA datasets, without any additional optimization. Specifically, the Social-IQA dataset is dedicated to commonsense question-answering (similar to the ECQA dataset), while the XCOPA dataset specializes in causal reasoning (similar to the COPA dataset). We adopt the Vicuna-7B as the targeted LLM, and Claude-2 as the explainer on these transfer tasks. The fidelity of the derived NL explanation on the target dataset is shown in Figure 6(b). The optimized trigger prompts show better explanation efficacy than human-crafted prompts when it is transferred from a dataset from a similar domain. This shows that the optimized trigger prompts generated by FaithLM possess a great property of data transferability.
@@ 4.4 Transferability of Optimized Trigger Prompt (RQ2)
(a) Evaluated by pre-trained NLI classifiers
(b) Evaluated by GPT-3.5
Figure 7: Ablation studies on evaluating contrary explanation. The results show that contrary explanations obtain opposite meanings to the derived explanations.
4.5 Ablation Studies on Contrary Explanation (RQ3)

### v4

@@ 3.2 FaithLM on Fidelity-enhanced Explanation
The framework of fidelity-enhanced explanation is illustrated in Figure 2(a). Since the initial explanation may be unreliable and unfaithful, we propose a fidelity-enhanced optimization approach designed to progressively generate explanations with higher fidelity. We aim to explain the response $\boldsymbol{Y}$ produced by the targeted LLM $f(\cdot)$ in response to the given input queries $\boldsymbol{X}$ with NL explanations $\mathcal{E}_{\text{NL}}$ following the goal of fidelity enhancement. In the first round of enhancement, the LLM explainer generates NL explanations $\mathcal{E}_{\text{NL}}$ following a given human-crafted explanation trigger prompt $\mathcal{P}_{\mathcal{E}}$ provided in Appendix I. The explanations are then generated by the explainer $\textsl{g}(\mathcal{P}_{\mathcal{E}}|~\boldsymbol{X},\boldsymbol{Y})$. Starting from the second round till converge, FaithLM collects a trajectory $\mathcal{T}$ with the NL explanations $\mathcal{E}_{\text{NL}}$ and their corresponding fidelity scores $\mathcal{S}_{\mathcal{E}}$ generated by Fidelity Evaluator. The collection process can be represented as $\mathcal{T}\leftarrow\{\mathcal{T},~[\mathcal{E}_{\text{NL}},~\mathcal{S}_{\mathcal{E}}]\}$, where $\mathcal{T}$ initially starts as an empty trajectory. Following this trajectory, the LLM explainer generates new explanations with the goal of achieving higher fidelity scores in subsequent iterations. This process is guided by the system prompts detailed in Figure 14.
@@
Algorithm of Fidelity-enhanced Explanation. The outline of FaithLM for Fidelity-enhanced Explanation is detailed in Algorithm 1. Specifically, in the first iteration, FaithLM generates the NL explanations using the human-craft prompts (line 1). starting from the second iteration till the convergence or optimization ends, FaithLM estimates the fidelity of the derived NL explanations (line 4). Then, we incorporate the explanation and its corresponding fidelity score to the trajectory (line 5), and update the explanations with the goal of achieving higher fidelity scores in subsequent iterations (line 6). The iteration terminates at a predetermined step or ceases earlier as soon as FaithLM observes a flipping performance from the targeted LLM $f(\cdot)$.
Algorithm 1 Fidelity-enhanced Explanation
@@
[…]
@@
3.3 FaithLM on Trigger Prompt Optimization
Despite the success of enhancing fidelity in Section 3.2, the low quality of the explanation trigger prompts $\mathcal{P}_{\mathcal{E}}$ may still hinder the optimization process of receiving a high-fidelity explanation. Given that the unknown preference for prompts from LLMs, human-crafted trigger prompts used in Fidelity-enhanced Explanation Optimization might lead to sub-optimal fidelity enhancement in the derived explanations. In this section, we hereby propose a new optimization pipeline under FaithLM, aiming to optimize the trigger prompt $\mathcal{P}_{\mathcal{E}}$ for generating NL explanations with higher fidelity scores as the LLM explanations of input each input query.
@@ 3.3 FaithLM on Trigger Prompt Optimization
The framework of Trigger Prompt Optimization is shown in Figure 2(b). The framework aims to optimize the trigger prompt to generate NL explanations with higher fidelity. Different from the optimization goal in Section 3.2, the trajectory in this task collects the trigger prompts $\mathcal{P}_{\mathcal{E}}$ and their fidelity scores $\mathcal{S}_{\mathcal{P}}$. The trajectory is constructed by the system optimization prompts detailed in Figure 13.
To estimate the fidelity score for a trigger prompt, FaithLM first adopts the randomly human-crafted trigger prompt to guide the LLM explainers to generate NL explanations, and then utilize the Fidelity Evaluator to assess the fidelity of the derived explanation. The final estimated score is averaged by the fidelity score $\mathcal{S}_{\mathcal{E}_{i}}$ of the hold-out dataset $(\boldsymbol{X}_{i},\boldsymbol{Y}_{i})\in\mathcal{D}$. Formally, the fidelity score for a trigger prompt $\mathcal{P}_{\mathcal{E}}$ is as follows:
@@
where $\mathcal{S}_{\mathcal{E}{i}}$ represents the fidelity score of the explanation $\mathcal{E}_{i}$, which is generated by $\textsl{g}(\mathcal{P}_{\mathcal{E}}|\boldsymbol{X}_{i},\boldsymbol{Y}_{i})$, as assessed by the Fidelity Evaluator.
@@
During the optimization, the trajectory begins from an empty set and starts to incorporate newly derived trigger prompts with the fidelity scores in each optimization iteration. Following this trajectory, the LLM explainer generates a new trigger prompt with the goal of achieving higher fidelity scores of explanations in subsequent iterations. After several rounds of iterations, FaithLM ultimately yields an optimal explanation trigger prompt with the highest fidelity score for the LLM explainer to generate a more faithful NL explanation.
@@
The outline of FaithLM for Trigger Prompt Optimization is detailed in Algorithm 2, which focuses on optimizing the trigger prompt for generating NL explanations. Specifically, in each iteration, LLM explainer $\textsl{g}(\cdot)$ leverages the trigger prompt to generate the NL explanations and estimates its fidelity (lines 4-7). The fidelity scores of the trigger prompts average the fidelity scores of the entire hold-out dataset (lines 8). Afterward, the trajectory appends the trigger prompt with its corresponding fidelity score (line 9), and updates the trigger prompt as a new sequence of words to achieve higher fidelity scores (line 10). Through multiple iterations, FaithLM progressively guides the trigger prompt to generate explanations with higher fidelity scores, following the reference path established in the trajectory. The iteration process terminates at a predetermined 20 step.
Figure 3: The fidelity evaluation of explanations on ECQA (left), TriviaQA-Long (middle), and COPA dataset (right). The reported scores are the average fidelity on testing instances in each step of fidelity-enhanced optimization.
@@ 4.1 Dataset and Baseline
Datasets. We evaluate FaithLM on three datasets with multiple tasks: ECQA Aggarwal et al. (2021) dataset on commonsense question-answer task, TrivaQA-Long Bai et al. (2023); Joshi et al. (2017) dataset on reading comprehension task, and COPA Kavumba et al. (2019); Roemmele et al. (2011) dataset on commonsense causal reasoning task. More details of datasets are provided in Appendix B. Baseline Methods. We compare FaithLM with two state-of-the-art baseline methods: SelfExp Madsen et al. (2024) and Self-consistency Wang et al. (2022). The former ones instruct LLMs to generate explanations using prompt engineering under single-forward inference, and the later ones leverage the outputs from the chain-of-thought prompting process as the model explanations.
Figure 4: Trustfulness evaluation of the NL explanations. Higher the proportion of “similar to ground-truth explanation," the more consistent the derived explanations are with the ground-truth NL explanations.
Comparison with Concurrent Faithfulness Notions.
Simulatability-based evaluations (e.g.,  Madsen et al. (2024); Chen et al. (2023a)) assess whether a model can reproduce predictions from an explanation, i.e., behavioral imitation. In contrast, FaithLM measures causal sensitivity via a controlled intervention on the explanation’s semantics: we compute $S_{E}=f(X)-f(X\,|\,\lnot E_{NL})$ and read off prediction shifts induced by a contrary hint, without token masking or task-specific heuristics. This aligns with intervention-driven faithfulness testing advocated by recent work and keeps the target model $f$ unchanged. [^1] See our definition and evaluator description in §3.1. Empirically, we additionally report a small-scale comparison where we compute (i) simulatability scores and (ii) $S_{E}$ on the same instances; we analyze where the two agree and where $S_{E}$ detects causal use beyond imitability.
4.2 Experimental Settings
We evaluate FaithLM through two complementary tasks: (i) fidelity-enhanced explanation generation and (ii) explanation-trigger prompt optimization. Details are presented in Appendix C.
Fidelity-Enhanced Explanation.
We measure the alignment between generated explanations and predictions Chen et al. (2025). For each example, FaithLM produces one NL explanation, and the average fidelity score across instances is reported as the final metric.
Explanation-Trigger Prompt Optimization.
This task optimizes the trigger prompt that guides FaithLM to generate more faithful explanations. At each iteration, 30 samples are drawn from the training data as a validation set, and the mean fidelity score on these samples is used as the optimization objective.
Evaluation Metrics. We evaluate explanation quality using two metrics: fidelity and truthfulness. Fidelity follows the intervention-based protocol of Chen et al. (2025), where contrary hints act as controlled semantic interventions and fidelity is the resulting prediction sensitivity. Truthfulness measures semantic consistency between generated and ground-truth explanations using GPT-4o, RoBERTa-Large, and XLNet-Large (Nie et al., 2020), following Liu et al. (2023). Evaluators classify each pair as similar, dissimilar, or non-relevant, and the proportion of “similar” outputs forms the truthfulness score. Full prompts appear in Appendix H.
Implementation Details. We use Vicuna-7B (Chiang et al., 2023) and Phi-2 (Javaheripi and Bubeck, 2023) as target models $f(\cdot)$, and GPT-3.5-Turbo and Claude-2 (Anthropic, 2023) as explainers $g(\cdot)$. The same models generate contrary hints. Results are averaged over three runs with grid search on hyperparameters. Both Vicuna-7B and Phi-2 share identical decoding settings. Full configurations and infrastructure details are given in Appendices E and F.
@@ 4.3 Explanation Efficacy of FaithLM (RQ1)
We assess the efficacy of derived explanations under the fidelity metric. FaithLM adopts the trajectory system prompts in Figure 14 of Appendix J. The generation of contrary hints is guided by the prompt in Table 8.
@@
(a) Evaluated by GPT-4o Evaluator
Figure 7: Ablation studies on evaluating contrary hint.
  • Fidelity Evaluation. The results in Figure 3 demonstrate that FaithLM achieves significantly higher fidelity scores across all three datasets compared with two baselines after 20 steps of optimization. Moreover, the optimization curve of fidelity demonstrates that 20 rounds of optimization are sufficient to converge. A similar phenomenon occurs across different settings of explainers and targeted LLMs. Additional results are provided in Appendix G.
  • Truthfulness Evaluation. To evaluate the truthfulness of explanations, we show the proportions of “similar to ground-truth explanations" in the ECQA dataset, as depicted in Figure 4. We leverage GPT evaluators and well-trained NLI evaluators to assess whether the given explanations are within similar content to ground-truth explanations. The results show that FaithLM’s explanations are more consistent with the ground-truth NL explanations, indicated by a larger proportion of “similar to ground-truth explanations" generated by FaithLM than baseline methods.
Efficacy of Explanation Trigger Prompts. We first show prompt optimization curves on three different datasets, and then leverage the optimal explanation trigger prompts to generate explanations via FaithLM. In the experiments, we randomly select 15 instances from the training dataset in each optimization round, and compute the average fidelity scores of the newly derived trigger prompts. After the progress is terminated, we evaluate the optimized trigger prompts on the testing set. The optimization step is uniformly established at 50 rounds across different explainer and targeted LLMs.
  • Trigger Prompt Optimization Curve. Figure 5 demonstrates the optimization curves of three datasets. We display the explainer as GPT-3.5-Turbo and Claude-2 and the explainer as Vicuna-7B. We observe that the optimization curve exhibits a generally ascending trend as the step progresses, interspersed with multiple waves throughout the optimization procedure. This indicates that FaithLM generates better explanation trigger prompts after the optimization. More results of optimization curves on remaining datasets are provided in Appendix G.
  • Explanation Generation by Optimized Trigger Prompts. We utilize the optimized explanation trigger prompts to generate explanations following Algorithm 1. The results are displayed in Figure 6(a), including the experiments conducted using all three datasets with Claude-2 as the explainer and Vicuna-7B as the targeted LLM. We observe that optimized explanation trigger prompts obtain higher fidelity scores than the initial human-crafted trigger prompt in generating explanations. This trend is consistent across all datasets, regardless of whether the explanations are refined by Algorithm 1.
4.4 Transferability of Trigger Prompt (RQ2)
We assess the transferability of ultimately optimized trigger prompts across different unseen datasets within the same domain, as depicted in Figure 6(b). Specifically, we transfer the optimized trigger prompts from the ECQA to the Social-IQA dataset, and from the COPA to the XCOPA datasets, without any additional optimization. Specifically, the Social-IQA dataset is dedicated to commonsense question-answering (similar to the ECQA dataset), while the XCOPA dataset specializes in causal reasoning (similar to the COPA dataset). We adopt the Vicuna-7B as the targeted LLM, and Claude-2 as the explainer on these transfer tasks. The fidelity of the derived NL explanation on the target dataset is shown in Figure 6(b). The optimized trigger prompts show better explanation efficacy than human-crafted prompts when it is transferred in similar domain. This shows that the optimized trigger prompts generated by FaithLM possess a great data transferability.
4.5 Ablation Studies on contrary hint (RQ3)